# Why Johnny Signs with Sigstore: Examining Tooling as a Factor in Software Signing Adoption in the Sigstore Ecosystem

### v3

@@ Abstract
Software signing is the strongest method for ensuring the integrity and authenticity of components in a software supply chain. Although software signing has been recommended by regulation and industry consortia, practical adoption of software signing is generally limited. Previous studies have recognized tooling as a key factor in influencing software signing adoption and quality. However, most research has primarily focused on its user interface aspects, with little attention given to other usability considerations, such as case-specific user challenges and factors influencing tool adoption and changes.
To understand how software tools influence the practice and adoption of software signing, we study the formative usability of Sigstore, a modern and widely adopted software signing tool. We interviewed thirteen (13) experienced security practitioners to study the factors that influence the selection of a tool, the problems associated with the use of such tools, how practitioners’ software signing tools have evolved, and what drives this migration. To summarize our findings: (1) We highlight the various factors practitioner’s consider before adopting a software signing tool; (2) We highlight the problems and advantages associated with the current tooling choices of practitioners; and (3) We describe the evolution of tooling adoption of our sample population. Our findings provide the software signing tool development community with valuable insights to improve their design of software signing tools.
1 Introduction
The reuse of software components in modern software products results in the formation of complex software supply chains, leading to an increased attack platform for these products [39, 79, 67]. A key security strategy to mitigate this expanding attack surface is ensuring the authenticity of actors and integrity of artifacts by establishing provenance [54, 68]. Software signing is the strongest guarantee for establishing provenance [69] and has been extensively advocated by academia, industry consortia, and government regulations [81, 16, 14]. To effectively implement this process, software signing tools–such as OpenPGP [55], Sigstore [53], and Notary [57]–have been proposed and are currently used in practice to sign software artifacts.
Although tooling has significantly advanced various technological fields [71, 41, 3, 47], its impact on software signing must be carefully examined to maximize its benefits. While prior "Why Johnny Can’t Encrypt" inspired studies have explored the usability of signing tools, their focus has been limited primarily to user interface design [89, 73, 63]. Furthermore, their evaluation techniques rely primarily on summative usability assessments (ease of use), which have been criticized for being difficult to measure and for providing limited insights to inform tool enhancements [13, 48]. The applicability of these studies is also constrained by their narrow focus on message (and email) signing and their reliance on novice volunteer participants, which limits the real-world relevance of their findings for software signing.
@@ 1 Introduction
Our work provides the first examination of the usability of tooling as a key factor influencing the adoption and practice of software signing. We conducted interviews with 13 experienced security practitioners to explore current software signing tools, with a focus on Sigstore as a case study. We apply a formative usability framework  [17], which emphasizes the iterative refinement of the studied tooling option, to develop our study instruments. Our study explores usability in two categories: users’ experience with the specified tooling option and the usability of the tooling options with respect to other software signing tools. Our study thus explores three key themes: the evolution and motivations behind software signing tool adoption, limitations and challenges encountered with currently adopted tooling, and the specific implementations of adopted tools.
@@
Our results highlight the factors practitioners consider when adopting specific software signing tools, using a case study of a key software signing tool. This serves as formative feedback for tool designers, offering insights into the drivers of tool usage. We present these findings through a clear articulation of practitioners’ challenges and the merits of using particular tools. Additionally, we found that practitioners’ decisions to adopt a software signing tool are influenced by less commonly discussed factors, such as the tool’s community, relevant regulations and standards, and certain functionalities of modern Software signing tools.
@@
    We report on the difficulties and advantages of using Sigstore, a next-generation software signing tool (section 5.1.1, section 5.1.2).
@@
    We discuss how practitioners’ software signing tools evolve over time and the factors that influence practitioners’ decisions to switch tooling options (section 5.2).
2 Background & Related Works
In this section, we discuss the foundations of software signing (section 2.1), usability studies in software signing (section 2.2), and the Sigstore ecosystem where we situate our work (section 2.3).
2.1 Software Signing
2.1.1 Signing Overview and Workflow
Software signing is a process that ensures the authorship and integrity of a software artifact by linking a cryptographic signature created by the software maintainer with their software artifact. Fundamentally, software signing is based on the public-key cryptography model  [69], and a traditional signing workflow follows a three-stage process of creation, distribution, and verification. We describe a typical signing workflow in Appendix C.
While the fundamental software signing process remains consistent across implementations, different signing tools and policies are shaped by the target platform, software artifacts, and root of trust. The choice of signing tools and the policies under which developers operate often depend on the operating system and distribution vendor. For example, Microsoft Authenticode [30] and Apple Xcode [33] provide platform-specific signing mechanisms, while package registries like PyPi [45] and Maven [1] enforce signing policies through dedicated tooling plugins. Developers also have a variety of software signing tools available to them, such as GNU Privacy Guard (GPG) [58], Signify [85], Sigstore [53], and Notary [57]. Each of these tools present different advantages and challenges, shaping how developers integrate signing into their workflows.
2.1.2 History and State-of-Art Tooling
Traditional Signing Tools: Many traditional signing tools, such as Git Commit Signing and GNU Privacy Guard are based on PGP. PGP (Pretty Good Privacy) is a proprietary encryption software originally developed by Phil Zimmerman in 1991. It is a public key cryptography-based protocol whose capabilities include encrypting files, creating and managing public and private keys, and creating digital signatures. Its open-source standard is OpenPGP [23, 55]. However, two main struggles associated with PGP-based software signing involve key management and establishing trust in public keys.
PGP relies on a "web of trust" model, where a signer’s public key is considered trustworthy based on endorsements from other trusted signers [23]. Each created public key is stored in a "key certificate" that specifies the public key, the name and/or email address associated with the public key, its creation date, and others who have signed the key. The signer is also responsible for secure management of their private key, preventing exposure, as the combination of their private key and passphrase enables the signing of artifacts [45].
State-of-art Signing Tools: Although software signing is widely recognized as an effective technique for mitigating software supply chain attacks, several obstacles hinder its adoption. Research highlights the prevalence of unsigned artifacts in software registries, emphasizing the need for improved signing practices [53, 90, 46]. This has led to the emergence of the next generation of software signing. Newman et al. introduced a new software signing platform in 2021 called Sigstore that addresses key adoption barriers, particularly key management, by leveraging keyless/ephemeral software signing techniques [53].
Keyless signing mitigates security risks associated with private key compromise. Instead of relying on a long-term private key, the signer uses a short-lived key that exists only for the duration of the signing process. This approach enhances security by preventing private key leaks while improving accessibility and usability [2]. As a result, many software registries are transitioning to this model to streamline verification and enforce stronger software integrity measures [15, 34]. While Sigstore’s development aims to solve the usability issues associated with traditional signing methodology, firsthand perspectives from practitioners using Sigstore will provide valuable insight into its actual usability and convenience.
2.2 Usability Studies in Software Signing
2.2.1 Usable Tools: Definition and Theory
Usable Tools: In basic terms, a software engineering tool assists developers in their software production process [27]. For software signing specifically, these tools are used to provide a method for which users can verify the integrity and authenticity of the code to be published, and developers use these tools to verify signatures of the software artifacts they may incorporate into their own projects. Deciding which software engineering tools to use is a critical question developers face when many options are available. A specific tool can affect a developer’s workflow efficiency, productivity [44], and the resulting quality of the work that is produced. This allows members of the computing industry to remain competitive by selecting the most suitable tools and environments that are conducive for the efficient production of software [27]. Additionally, tooling is important because it enables engineers to use software development practices they have not in the past [11].
Usability Evaluation: Studies assessing implementation strategies of products are typically framed as usability evaluations [49]. Usability evaluations are generally categorized into two forms: summative (validation or ease of use) and formative (diagnostic usability) [48]. Summative usability refers to the extent to which a system or product can be used to achieve a specified goal effectively, efficiently, and satisfactorily within a given context [35, 65]. In contrast, formative usability focuses on product assessments aimed at identifying areas for improvement [82]. Although summative usability evaluations are more commonly used, they are not designed to provide actionable feedback for enhancing the product [48]. Summative evaluations primarily measure metrics such as effectiveness, efficiency, and user satisfaction [35]. However, such metrics have been widely criticized for being subjective and difficult to quantify [48, 13].
In contrast, formative usability evaluations aim to diagnose user challenges and provide iterative feedback to enhance product functionality [82, 48]. These evaluations emphasize real-world usability issues and are more suited for refining a product’s design. While summative evaluations are more common, their limitations highlight the need for formative assessments when improving software signing tools.
2.2.2 Empirical Studies
Signing Tooling Usability Studies: The usability of digital signing tools has been widely studied, though often within the context of encryption rather than broader software signing applications. Whitten and Tygar’s seminal study, Why Johnny Can’t Encrypt [89], found that PGP 5.0’s user interface was too complex for non-expert users, making encryption and signing inaccessible. Sheng et al. [73] later observed that PGP 9’s interface had not significantly improved, particularly in key verification, encryption transparency, and digital signing usability. Other studies [10, 64, 60] have highlighted persistent usability issues, including the complexity of the public-key model.
Further research has explored the differences between manual and automated signing tools. Fahl et al. [19] and Ruoti et al. [62] found no significant usability differences between these approaches. Meanwhile, Atwater et al. [5] demonstrated that users preferred automated solutions based on the PGP standard. While these studies provide insights into encryption tools, their relevance to broader software signing workflows is limited.
@@ 2.2.2 Empirical Studies
These studies exhibit weaknesses in their scope, methodology, and the relevance of their findings to software signing. First, they primarily focus on evaluating the usability of signing tooling in the context of email or message encryption, limiting their application to software signing. Additionally, they limit their scope to user interface design, overlooking critical factors such as tooling adaptability and macroenvironment (e.g., tooling user communities). Furthermore, much of this research is outdated, reflecting decades-old technologies and failing to incorporate contemporary practitioner experiences. Their methodology predominantly employs summative usability assessments, which focus on ease of use [13]. These studies typically rely on novice volunteers to evaluate usability. In addition to the flaws of the summative evaluation approaches, the reliance on novice participants limits the applicability of their findings. Evaluating usability through inexperienced users does not yield specialized feedback that can effectively guide tool enhancements [42, 50].
2.2.3 Related Software Signing Research
The use of software signing as a security method to mitigate Software Supply Chain security has been studied in the literature. These studies has focussed on the adoption and practice of Software signing. Schorlemmer et al. [69] examine factors influencing software signing adoption through a framework of signing events. They identify tooling as a major factor affecting the quality of software signatures but focus primarily on the role of dedicated tooling within software registries, rather than its broader impact on engineering workflows. In our parent study, Kalu et al. [37], we provided valuable insights into the implementation and adoption frameworks of software signing, with a focus on broad software signing challenges faced in commercial settings. These challenges included issues of usability, compatibility, and integration. However, while we highlighted these broad challenges, our findings lacked the specificity needed to identify which tools are most impacted by these issues and under what conditions of use, limiting practitioners’ contextual understanding of these problems.
These prior studies acknowledge that improved tooling can drive the adoption of software signing. However, they fail to clearly define what constitutes an "improved tool" from the perspective of practitioners. While they emphasize broad usability factors, they overlook the specific technical and workflow-related requirements that practitioners consider when selecting tools for real-world use.
2.3 Sigstore
Due to the emergence of next-generation software signing and the growing adoption of Sigstore as a signing platform, we focus on the Sigstore tool specifically in our research.
Importance: Sigstore, launched in 2021, was developed to address critical challenges in software signing, including complex key management, usability barriers, and software supply chain security risks. Traditional signing tools, such as OpenPGP and enterprise certificate authorities, rely on long-term cryptographic key storage, which introduces risks of key compromise, revocation difficulties, and usability friction. In contrast, Sigstore simplifies the signing process by providing keyless, ephemeral authentication mechanisms. Within 13 months of its public release, Sigstore recorded 46 million artifact signatures  [29]. It has since gained significant adoption and support from open-source ecosystems  [84, 25, 31, 6, 86] and various organizations  [38, 76, 75, 77]. Given its growing adoption and integration in major software ecosystems, Sigstore represents a state-of-the-art software signing approach, making it a compelling case study for examining usability and security trade-offs.
Design Aspects Affecting Usability: One key challenge that is associated with traditional signing methods is key management, which is addressed by Sigstore. In conventional certificate authority-based signing, such as Windows Authenticode, a software artifact’s validity is linked to the certificate’s status—whether it is still valid, expired, or revoked [43, 70]. A revoked certificate can indicate compromise, but it may not be immediately clear when the compromise occurred [40]. This uncertainty allows some revoked certificates to be used for signing malware while still appearing valid to verification systems. Sigstore mitigates this risk by issuing short-lived certificates. These certificates significantly reduce the chances of key and certificate compromise since they automatically expire after a short duration, limiting their potential misuse.
Complexity: Sigstore uses a modular design of several components to implement software signing. As described by [53], Sigstore consists of eight key components: signers, verifiers, the artifact log, an OpenID Connect (OIDC) provider, a certificate authority, monitors, transparency logs, and package repositories. While Sigstore is designed to function as a cohesive system, these components can also be implemented individually based on specific use cases. The signing process starts with the signer using OpenID Connect (OIDC) to obtain a signing certificate. Unlike traditional methods requiring long-term key management, Sigstore associates a signer’s identity with short-lived tokens issued by services such as Google, GitHub, and Microsoft. Next, a software signing certificate is issued to the signer based on their verified identity. In this step, Fulcio serves as the certificate authority (CA), ensuring that the OIDC token was correctly issued to the requesting signer before granting the certificate. To monitor signing behavior, a verifier makes use of Fulcio’s log and the transparency log (Rekor) to verify for identity of the creator of the signatures, the certificate validity, and signature validity.
3 Knowledge Gaps & Research Questions
The existing literature on software signing, as outlined in  section 2, has examined its practice and adoption across different contexts, including private and open-source ecosystems. These studies highlight the critical role that the choice of a software signing tool plays in ensuring the quality of the software signing process.
Although certain aspects of software signing tools have been explored, such as the usability of the user interface design and functional exclusivity (dedication of tooling resources), the broader impact and role of software signing tools remain largely understudied. The findings of these works underscore that tooling is a multi-faceted and multi-layered factor. A general usability evaluation of these tooling options would significantly enhance their adaptability and facilitate improvements. Also, the existing research on software signing tooling primarily relies on experiments with novice users and quasi-experiments to investigate isolated aspects of tooling. Practitioner perspectives on the nuanced role of tooling in shaping the use and practice of software signing remain limited, particularly regarding how the usability of these tooling contributes to their adoption.
Usability Framework:
Figure 1: Usability Framework from Cresswell et al. [17]. Their framework assesses usability based on four factors: Technological (T), Social/Human (People), Organizational (O), and Macroenvironmental (M) factors. Our framework applies their framework to two concerns of software signing tool usage: (1) the tool itself and (2) previous signing experiences.
To examine the usability of software signing tooling, an examination of experienced practitioner experiences is necessary. As noted in section 2.2, the summative usability approach mostly is mostly hindered by its use of subjective measurement metrics. To effectively, answer usability in terms of how software signing tooling are used in practice, a formative usability framework is warranted. A formative framework addresses the experiences of regular users of a tool, rather than controlled experience of experimental subjects  [48].
Our formative usability evaluation framework, Figure 1, is based on Cresswell et al.’s [17] formative framework. Their framework identifies four key factors that influence the overall usability of a product or technology: Technological (T), Social/Human (People), Organizational (O), and Macroenvironmental (M) factors. Usability is determined by the interplay of these factors.
Our formative evaluation applies of their four-factor framework to two key areas: users’ experiences with the specified tooling option and the usability of the tooling option compared to previous software signing tools. We designed our study instrument to address these broad usability themes, emphasizing the identified usability factors.
Research Questions: Our research addresses the following questions in light of the positions outlined above within the context of Sigstore:
@@ 3 Knowledge Gaps & Research Questions
    What drawbacks and benefits of software signing tools influence their adoption in practice?
@@
    What factors influence practitioners’ decisions to switch software signing tools over time?
4 Methodology
Figure 2: Study Methodology. We developed our interview protocol by reviewing Sigstore grey literature and the academic literature using our adopted formative usability framework. Finally, we analyzed our results using a thematic analysis informed by our usability evaluation framework (Figure 1).
We undertook a qualitative case study of the Sigstore software signing ecosystem through interviews of engineers. We applied a formative usability framework to identify the key factors that impact the adoption of Sigstore in practice.
Our methodology is illustrated in Figure 2. This study was approved by our Institutional Review Board (IRB).
4.1 Research Design and Rationale
4.1.1 Rationale

### v4

# leftmargin=0.15cm, rightmargin=0.15cm, font=italic, vskip=0.10cm, indentfirst=false

@@ Abstract
Software signing is the most robust method for ensuring the integrity and authenticity of components in a software supply chain. However, traditional signing tools place key management and signer identification burdens on practitioners, leading to both security vulnerabilities and usability challenges. Next-generation signing tools such as Sigstore have automated some of these concerns, but little is known about their usability and adoption dynamics. This knowledge gap hampers the integration of signing into the software engineering process.
To fill this gap, we conducted a usability study of Sigstore, a pioneering and widely adopted exemplar in this space. Through 18 interviews, we explored (1) the factors practitioners consider when selecting a signing tool, (2) the problems and advantages associated with the tooling choices of practitioners, and (3) practitioners’ signing-tool usage has evolved over time. Our findings illuminate the usability factors of next-generation signing tools and yield recommendations for toolmakers, including: (1) enhance integration flexibility through officially supported plugins and APIs, and (2) balance transparency with privacy by offering configurable logging options for enterprise use.
I Introduction
The reuse of software components in modern products creates complex supply chains and expands the attack surface for these systems [1, 2, 3]. A core strategy for mitigating this risk is to establish provenance—verifying actor authenticity and artifact integrity—through software signing [4, 5]. Software signing is the strongest guarantee of provenance [6] and is widely advocated by academia, industry consortia, and government regulations [7, 8, 9]. High-profile supply-chain attacks—such as the SolarWinds breach—have further underscored the need for stronger, more automated signing infrastructure. In response, tools have evolved into two classes: traditional solutions (e.g., OpenPGP [10]), which delegate key management and signer identification to practitioners, and next-generation, identity-based tools (e.g., Sigstore [11]), which further automate the process via short-lived keys and external identity providers.
Tooling has significantly advanced various technological fields [12, 13, 14, 15] and has often played a key role in the adoption of these technologies [16, 17]. There is a continual need to evaluate the adoption practices and usability of such tools to inform practitioners, potential users, researchers, and toolmakers of their effectiveness in different contexts [18]. Traditional software signing tools have been studied in this way, most notably in several “Why Johnny Can’t Encrypt” inspired studies[19, 20, 21]. Although next-generation signing tools promise greater automation, there remains a lack of comprehensive usability and adoption analyses for these solutions. Such studies could assess how automation affects adoption suitability, elicit practitioners’ desired improvements, and guide future tool design and research.
@@ I Introduction
Our work offers the first examination of next-generation software signing tools’ usability as a critical factor influencing their adoption and practice. Building on and improving previous methodologies — which relied on novice users and summative usability assessments criticized for their lack of actionable feedback[18, 22] — we conducted interviews with 18 experienced security practitioners to study current software signing tools, focusing on Sigstore as a case study. We analyzed our data with a formative usability framework [23], focusing on two usability dimensions: practitioners’ experiences with Sigstore and its usability relative to other signing tools. From these interviews, we distilled three key themes: the evolution and motivations behind tool adoption, the limitations and challenges of currently adopted tools, and the specific implementation practices shaping real-world signing workflows.
@@
Our results highlight the usability factors practitioners consider when adopting next-generation software signing tools. This serves as formative feedback for tool designers and researchers, offering insights into the drivers of tool usage. We present these findings through a clear articulation of practitioners’ challenges and the merits of using particular tools. Additionally, we found that practitioners’ decisions to adopt a software signing tool are influenced by less commonly discussed factors, such as the tool’s community, relevant regulations and standards, and certain functionalities of next-generation Software signing tools.
@@
    We report on the difficulties and advantages of using Sigstore, a next-generation software signing tool.
@@
    We discuss how practitioners’ choices of software signing tools change over time and the factors influencing those decisions.
Significance of Study: Our work targets Sigstore—a pioneer next-generation signing tool that has seen rapid uptake across major open-source ecosystems[24, 25, 26], cloud-native projects, and corporations[27, 28, 29, 30, 31], involving a large number of engineers worldwide. By unpacking the usability factors that drive or hinder Sigstore’s adoption, we not only inform the ongoing refinement of its automation workflows (e.g., certificate issuance and key management) but also provide a template for evaluating and improving similar Next-generation(identity-based) signing solutions. Also, since Sigstore exemplifies a broader class of tools that embed security into automated build and deployment processes, our findings can be potentially generalized to guide the design and implementation of other automated software-supply-chain technologies, ultimately advancing the usability—and thus the security—of automated software engineering practices.
II Background & Related Works
In this section, we review software signing technologies (section II-A) and usability assessments (section II-B).
II-A Software Signing
Software signing ensures the authorship and integrity of a software artifact by linking a maintainer’s cryptographic signature to their software artifact. The implementation of software signing is shaped by several factors: the software artifacts to be signed (e.g., driver signing[32]), the root-of-trust design (e.g., public key infrastructure/PKI [33]), and ecosystem policies (e.g., PyPi[34] and Maven[35]).
II-A1 Software Signing Tools
Software signing tools can be classified into two groups: traditional and next-generation (identity-based) [5]. Traditional signing tools such as GPG [36] follow a conventional public-key cryptography approach, relying on long-lived asymmetric key pairs and delegating key-management responsibilities to the signer. Trust is established through endorsements by other signers.
Next-generation (identity-based) signing tools automate key management by issuing short-lived, ephemeral certificates tied to a signer’s identity via external OpenID Connect (OIDC) or OAuth providers. This “keyless signing” approach — exemplified by Sigstore [11], OpenPubKey [37], and SignServer [38] — eliminates long-term key handling and embeds transparency through append-only logs, improving usability and security. We discuss the architecture and workflow for next-generation signing tools in section IV-B using the case of Sigstore.
II-A2 Effects of Signing Tools on Software Signing Adoption
Previous studies recognize that better tooling influences software-signing adoption. For example, Schorlemmer et al. [6] performed a measurement study of several software package registries, identifying tooling as a key determinant of signature quality. Similarly, Kalu et al. [39] explore implementation and adoption frameworks in commercial settings, highlighting tool usability, compatibility, and integration challenges. While these works emphasize broad usability factors, they offer little detailed guidance on what makes a signing tool usable, overlooking the fine-grained technical and workflow requirements that drive real-world tool adoption.
II-B Usability Studies in Software Signing
II-B1 Usability – Definition, Theory & Evaluation Frameworks
Tools enable engineers to adopt new software development practices [40]. A usable tool is one whose functionality effectively supports its intended purpose—achieving product goals [18, 41]. Studies assessing implementation strategies of tools are typically framed as usability evaluations [42].
To evaluate a tool’s usability, studies typically follow a usability evaluation framework [42]. There are two kinds: summative, which measures a tool’s effectiveness [43, 44], and formative, which provides iterative feedback for improvement [18, 45]. Given the aims of this work, we took a formative approach. We considered several formative frameworks [46, 47, 48, 49] and found that Cresswell et al.’s four-factor framework [23] helped explain our data the best. That framework is summarized in Figure 1.
Figure 1: Cresswell’s Usability Framework [23]. It assesses usability based on four factors. Technological (T) concerns the technical characteristics of the system (e.g., performance, ease of use) and data management (availability, integrity). Social/Human (P) relates to users and their interactions with the tool, e.g., user satisfaction, engagement, prior experience, and correct use of features. Organizational Context (O) are factors internal to participants’ organization, e.g., training and leadership support. Macroenvironmental (M) are external influences, e.g., regulations, economics, and vendor relations.
II-B2 Empirical Studies
The usability of traditional signing tools—often studied in the context of encryption—remains a significant challenge. Whitten and Tygar’s seminal study, “Why Johnny Can’t Encrypt” [19], found PGP 5.0’s interface too complex for non-experts, a problem that persisted in PGP 9 [20], especially around key verification, transparency, and signing. Subsequent work [50, 51, 52] has underscored the public-key model’s inherent complexity. Research comparing manual versus automated signing tools has yielded mixed results: some works observed no clear usability advantage [53, 54], whereas Atwater et al. showed a user preference for automated PGP-based solutions [55]. Although informative, these studies focus on traditional encryption scenarios (e.g., email and messaging), limiting their applicability to broader software-signing workflows.
@@ II-B2 Empirical Studies
These studies exhibit weaknesses in their scope, methodology, and the relevance of their findings to software signing. First, they primarily focus on evaluating the usability of signing tooling in the context of signing communication (e.g., email, messages), limiting their application to signing software (i.e., securing the software supply chain). Furthermore, much of this research is outdated, reflecting decades-old technologies and failing to incorporate contemporary practitioner experiences. Their methodology predominantly employs summative usability assessments that rely on novice volunteers, which limits the applicability of their findings for guiding tool enhancements [56, 57].
III Knowledge Gaps & Research Questions
Existing literature on software signing (see section II) has examined practice and adoption across private and open-source ecosystems, highlighting the importance of tool choice in signature quality. Prior usability studies have focused exclusively on traditional signing tools and have been limited to evaluating user-interface design. However, next-generation signing tools remain understudied, creating a need for:
  • A holistic evaluation of how software signing tools influences software-signing adoption and practice, and how usability affects this process.
  • Practitioner perspectives on the nuanced role of tooling in shaping software-signing use and practice. Current research (see section II-B2) is limited because it primarily relies on experiments with novice users, while studies in section II-A2 use quasi-experiments or general practitioner interviews without a specific focus on signing tools.
In light of these gaps, a comprehensive usability evaluation of next-generation signing tools would offer valuable insights into their role in adoption, improve their design, and enhance their adaptability.
Research Questions: Building on identified gaps in next-generation tooling usability and its impact on adoption, we frame our research around Sigstore—our chosen next-generation signing tool case study—and ask:
@@ III Knowledge Gaps & Research Questions
    What strengths and weaknesses of Sigstore influence its adoption in practice?
@@
    What factors influence practitioners’ decisions to change the software signing tools they use?
IV Methodology
We selected a case study method to answer our RQs. Our methodology is illustrated in Figure 2. This section provides our design rationale (section IV-A), defines the case study context (section IV-B), discusses instrument design (section IV-C), data collection (section IV-D), and analysis (section IV-E).
Our Institutional Review Board (IRB) approved this work.
IV-A Rationale for Case Study